OpenStudy (anonymous):

what is the PH of a solution containing .03 M of acetic acid and .3 mole of sodium acetate (CH3COONa) ? Ka for CH3COOH = 1.8*10^-5

OpenStudy (anonymous):

CH3COONa ----> CH3COO- + Na+ CH3COONa dissociates completely producing 0.3 moles of CH3COO- ion in 1 litre of the solution.So initially the concentration of CH3COO- ion becomes 0.3M CH3COOH ---> CH3COO- + H+ 0.3M 0.3M CH3COOH ---> CH3COO- + H+ .... Ka=1.8*10^-5 0.3M-x 0.3M+x x \[([0.3M+x] [x])\div([0.3-x]])\]=Ka = 1.8*10^-5 x= concentration of H+ ions.Since Ka is very small, of the order of 10^-5,the acid dissociates very poorly and x is negligible compared to 0.3.So 0.3-x and 0.3 +x is approx. 0.3 .So those two terms cancel therefore x=1.8*10^-5 pH= -log(1.8*10^-5)

OpenStudy (aaronq):

for buffer systems you can also use the Henderson-Hasselbalch equation: \(pH=pKa+log\dfrac{[A^-]}{[HA]}\)